what are the roots of (x-76)^2 =y

Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)

y = (x-76)^2
(x-76)^2 = 0
(x-76) = 0
(x-76) = 0, x = 76

Answer by Alan3354(69443)